OpenStudy (anonymous):

if A+B+C=180 prove that.. cos^2a(b+c/2)+cos^2a(a/2)=1

hartnn (hartnn):

if a+b+c = 180 then whats (b+c)/2 ?

hartnn (hartnn):

in terms of a ?

hartnn (hartnn):

can u find ?

OpenStudy (anonymous):

What @hartnn wants you to realize is that: \[ \cos(90^\circ -x) = \sin(x) \]

OpenStudy (anonymous):

b+c/2=90-a/2

hartnn (hartnn):

correct, now whats cos^2 (90 - a/2) = .... ? use what wio said

OpenStudy (anonymous):

cos^2(b+c/2)=sin^2a/2

hartnn (hartnn):

now a/2 is an angle, like x so whats sin^2 x + cos^2 x =....?

OpenStudy (anonymous):

1

hartnn (hartnn):

right, so sin^2 (a/2) + cos^2 (a/2) will also be = 1 :)

OpenStudy (anonymous):

thts it ?

hartnn (hartnn):

thats what u wanted to prove, that expression =1 we simplified it and got till sin^2(a/2) + cos^2(a/2) which indeed =1

hartnn (hartnn):

so yeah

OpenStudy (anonymous):

thankss :)

hartnn (hartnn):

welcome ^_^